OpenStudy (anonymous):

I have z = 2cis(theta), what is -z in polar form?

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.

OpenStudy (anonymous):

@AllTehMaffs ... could you help me?

OpenStudy (anonymous):

yeah. Hi! So \[ z =2 cis \theta = 2 \cos \theta + 2i\sin \theta\] right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

factor out the 2 and it looks like \[z = 2(\cos \theta + i \sin \theta) \] Euler's formula is \[\cos \theta + i \sin \theta = e^{i\theta}\] and polar form looks like \[z = re^{i \theta} \] Soo.....

OpenStudy (anonymous):

now that's something i have not learn yet. @@ i only started the basics of complex numbers.

OpenStudy (anonymous):

You've not seen Euler's formula? hmmm... 1 sec... What have you learned about so far?

OpenStudy (anonymous):

yea.... i have not seen that formula. I have learnt the basics of it, only \[z=rcis \theta\] multiplying and dividing in polar form, conjugate. That's all

OpenStudy (anonymous):

okay, cool.. I'm just confused because becase 2cistheta *is* polar form.... so I don't know what they want you to convert it to >.< could it just be \[z = 2cis \theta \\ \ \\ -z = -2cis \theta \]? Do they want you to find theta?

OpenStudy (anonymous):

hmmm.... they said that it can't be \[-z = -2cis \theta\] cuz if its written in \[z= rcis \theta\] , r needs to be positive. The answer is \[z=2cis(\theta+\pi)\] but no idea how they got that

OpenStudy (anonymous):

OH! They are correct, and I totally forgot about that. \[ -z = -2cis \theta = -2 \cos \theta -2 \sin \theta\]  \[ \cos (\theta + \pi) = - \cos \theta \] \[ \sin( \theta + \pi) = - \sin \theta \] Those are just trig identities. You can see they're true if you look at the (very poorly drawn) picture.. but using those \[ \\ -z = 2 (-\cos \theta) +2 (-\sin \theta) \\ \ \ \ = -2 \cos (\theta + \pi) -2 \sin (\theta + \pi)\ \\ \ \ \ = 2 \cos (\theta + \pi) +2 \sin (\theta + \pi) \\ = 2 cis (\theta + \pi) \] does that make sense?